# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Now we assume that the block $b$ is a hyperfocal abelian Frobenius block. In the following, we will give a more explicit description of the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $X$ in the decomposition (4.6.1) under this assumption. To do this, we need to investigate the Brauer quotients of $\mathbb{A}_{\delta}$ at a class of special nontrivial subgroups of $D$.
@@
Denote $N_{P}(Q)$ by $P_{Q}$. It is clear that $b_{Q}$ is still a block of $P_{Q}C_{G}(Q)$ over $\mathcal{O}$. We denote it by $c_{Q}$ when the block $b_{Q}$ is viewed as a block of $P_{Q}C_{G}(Q)$. By assumption, $P_{Q}$ is a defect group of the block $c_{Q}$. Since the block $b_{Q}$ is nilpotent by Proposition 3.12, the block $c_{Q}$ of $P_{Q}C_{G}(Q)$ is also nilpotent. Since $P_{Q}$ contains $D$, there is a unique local pointe $\varepsilon$ of $P_{Q}$ on $\mbox{$\mathcal{O}$}Gb$ such that $(P_{Q})_{\varepsilon}\leq P_{\gamma}$ by [25, Proposition 4.2]. Now we can take an element $l$ of $\varepsilon$ such that $l$ is in $\mathbb{A}$ by [25, Proposition 2.8] and $lj=jl=j$.
By [29, Lemma (40.2)], $\mathrm{Br}_{Q}(l)$ belongs to a local point of $P_{Q}$ on $kC_{G}(Q)$. Since $P_{Q}$ is a defect group of the block $c_{Q}$, $\mathrm{Br}_{Q}(l)kP_{Q}C_{G}(Q)\mathrm{Br}_{Q}(l)$ is a source algebra of the block $\bar{c}_{Q}$ over $k$. Since the block $c_{Q}$ is nilpotent, by [17, Theorem 8.12.3], there is a an indecomposable endopermutation $kP_{Q}$-module $V_{P_{Q}}$ with vertex $P_{Q}$ such that, setting $\bar{T}_{Q}=\mathrm{End}_{k}(V_{P_{Q}})$, we have an isomorphism of $P_{Q}$-interior algebras
$\mathrm{Br}_{Q}(l)kP_{Q}C_{G}(Q)\mathrm{Br}_{Q}(l)\cong\bar{T}_{Q}\mathop{\otimes}\limits_{k}kP_{Q}$
@@
It is well-known that there is a unique local point,denoted by $\tilde{\delta}$, of $D$ on $\bar{T}_{Q}$. Take $f\in\tilde{\delta}$ and set $V_{Q}=fV_{P_{Q}}$. Then $V_{Q}$ is an indecomposable endopermutation $kD$-module with vertex $D$. Clearly, the action of $Q$ on $V_{Q}$ is also trivial. So $V_{Q}$ can be viewed as a $k(D/Q)$-module with vertex $D/Q$. Set $\bar{S}_{Q}=\mathrm{End}_{k}(V_{Q})\cong f\bar{T}_{Q}f$. By [17, Proposition 7.4.4], the $D$-interior algebras $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kC_{P}(Q)$ and $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ are both primitive. Note that $\mathrm{Br}_{Q}:\mathbb{A}^{D}\rightarrow\mathbb{A}(Q)^{D/Q}$ is surjective and $\mathrm{Br}_{D/Q}\circ\mathrm{Br}_{Q}$ is just the Brauer homomorphism $\mathrm{Br}_{D}$ from $\mathbb{A}^{D}$ to $\mathbb{A}(D)$. Then $\mathrm{Br}_{Q}(j)$ belongs to a local point of $D$ on $((l\mathbb{A}l)(Q))^{D}$. Comparing the isomorphism (4.10.3) with the isomorphism (4.10.4), we get $(lAl)(Q)=\mathop{\bigoplus}\limits_{uD\in C_{P}(Q)/D}(l\mathbb{A}l)(Q)\cdot u$. Therefore, by [25, Proposition 2.8], $\mathrm{Br}_{Q}(j)$ also belongs to a local point of $D$ on $((lAl)(Q))^{D}$. Note that by [22, 5.3], local points of $D$ on $(lAl)(Q)$ and $(l\mathbb{A}l)(Q)$ are both unique. Therefore, we can assume that the isomorphism (4.10.3) maps $\mathrm{Br}_{Q}(j)$ to $f\mathop{\otimes}\limits_{k}1_{C_{P}(Q)}$. Then the isomorphisms (4.10.1) and (4.10.2) are obtained.
Borrowing the notation from the paragraph 3.1, we have $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}j\mathbb{A}uj$ with $a_{u,\delta}\in A_{\delta}^{*}\cap\mathbb{A}u$ satisfying $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ for any $w\in D$. So $(j\mathbb{A}uj)(Q)=\mathbb{A}_{\delta}(Q)\cdot\mathrm{Br}_{Q}(a_{u,\delta})$ when $u$ is in $C_{P}(Q)$. In particular, $\mathrm{dim}_{k}(\mathbb{A}_{\delta}(Q))=\mathrm{dim}_{k}((j\mathbb{A}uj)(Q))$ when $u$ is in $C_{P}(Q)$. By isomorphisms (4.10.3) and (4.10.4), we have $\mathrm{dim}_{k}(A_{\delta}(Q))=\frac{|C_{P}(Q)|}{|D|}\mathrm{dim}_{k}(\mathbb{A}_{\delta}(Q))$. We are done. ∎
In general, the strucutre of $\mathbb{A}_{\delta}(Q)$ above holds for arbitrary nontrivial subgroup $Q$ of $D$. And this general structure will help us construct a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ in the next section.
@@
Moreover, if $V_{Q}^{\prime}$ is another indecomposable endopermutation $k(D/Q)$-module with vertex $D/Q$ such that setting $\bar{S}_{Q}^{\prime}=\mathrm{End}_{k}(V_{Q}^{\prime})$, $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras, then $V_{Q}$ is isomorphic to $V_{Q}^{\prime}$ as $kD/Q$-modules.
@@
It is well-known that there is some $x\in G$ such that ${{}^{x}}(Q,\bar{b}_{Q})\leq(P,\bar{b}_{P})$ and $C_{P}({{}^{x}}Q)$ is a defect group of the block ${{}^{x}}b_{Q}$ of $C_{G}({{}^{x}}Q)$ over $\mathcal{O}$. Equivalently, $C_{P^{x}}(Q)$ is a defect group of the block $b_{Q}$. Since $D$ is abelian, by [33, Theomre 2], we can assume that $x$ is in $N_{G}(D_{\delta})$. Furthermore, since $N_{G}(D_{\delta})/C_{G}(D)=PC_{G}(D)/C_{G}(D)\cdot E_{\mathfrak{h}}$, we can assume that $xC_{G}(D)$ is in $E_{\mathfrak{h}}$. Then we have $\mathfrak{a}_{x}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{x}(wj)\mathfrak{a}_{x}^{-1}=(xwx^{-1})j$ for any $w\in D$.
@@
Clearly, $A^{x}$ is also a source algebra of the block $b$ determined by the defect pointed group $P_{\gamma}^{x}$. Since $x$ is in $N_{G}(D_{\delta})$, we have $\mathbb{A}^{x}$ is a hyperfocal subalgebra of $A^{x}$ with hyperfocal subgroup $D$. The conjugation action of $x$ induces an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{{{}^{x}}Q}\longrightarrow(\mathbb{A}^{x})_{\delta^{x}}^{Q}$. At the same time, the conjugation action of $\mathfrak{a}_{x}$ induces an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{Q}\longrightarrow\mathbb{A}_{\delta}^{{{}^{x}}Q}$. Composing these two isomorphisms together, we get an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{Q}\longrightarrow(\mathbb{A}^{x})_{\delta^{x}}^{Q}$, which is in fact a $D$-interior algebra isomorphism. By Lemma 4.10, we get this general structure of $\mathbb{A}_{\delta}(Q)$.
@@
Let $V_{Q}^{\prime}$ be another indecomposable endopermutation $k(D/Q)$-module with vertex $D/Q$ fulfilling the property in this lemma. Then we have $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Tensoring with $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}$- on the both sides of this isomorphism yields a $D$-interior algebra isomorphism $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$. By [22, 5.3], local points of $D$ on $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime}$ and $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ are both unique. We take two elements $e^{\prime}$ and $e$ belonging to these two unqiue local points, respectively. Then by [17, Theorem 7.4.2], $e(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD)e$ is isomorphic to $kD$ as $D$-interior algebras. It is clear that $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}$ is also a primitive Dade $D$-interior algebra. By [17, Proposition 7.4.4], we have $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}\mathop{\otimes}\limits_{k}kD$ is a primitive $D$-interior algebra. Hence, we have $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}\mathop{\otimes}\limits_{k}kD\cong kD$ as $D$-interior algebras. This implies that $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}$ is a trivial $D$-interior algebra. Then the elements in the Dade group of $D$ corresponding to $\bar{S}_{Q}$ and $\bar{S}_{Q}^{\prime}$ are inverses of each other. In particular, $V_{Q}\cong V_{Q}^{\prime}$ as $k(D/Q)$-modules. ∎
@@
Now we can refine the $\mbox{$\mathcal{O}$}(D\times D)$-structure of $\mathbb{A}_{\delta}$ in the case where the block $b$ is a hyperfocal abelian Frobenius block.

### v2

@@ 4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Let $M$ be an $\mathbb{A}$-module. Fix an element $u$ in $P$. We can define the twisted $\mathbb{A}$-module ${{}_{u}}M$ by setting ${{}_{u}}M=M$ as an $\mathcal{O}$-module, with $u\mathfrak{a}u^{-1}$ acting on $m\in{{}_{u}}M$ as $\mathfrak{a}$ on $m\in M$ for any $\mathfrak{a}\in\mathbb{A}$ and any $m\in M$. Furthermore, recall from the paragraph 3.1 that there is an element $a_{u,\delta}$ in $A_{\delta}^{\times}$ with $\mathbb{A}_{\delta}^{a_{u,\delta}}=\mathbb{A}_{\delta}$. Then we can similarly define the twisted $\mathbb{A}_{\delta}$-module ${{}_{u}}M_{\delta}$ for any $\mathbb{A}_{\delta}$-module $M_{\delta}$ through the conjugation action of the element $a_{u,\delta}$. Since the choice of the element $a_{u,\delta}$ is unique up to a multiplication by an element in $(\mathbb{A}_{\delta}^{D})^{\times}$, so the twisted module ${{}_{u}}M_{\delta}$ is independent of the choice of the element $a_{u,\delta}$ up to isomorphism. At the same time, letting $C$ be a $P$-algebra, we have the analogous notation ${{}_{(u,u)}}{\bf M}$ and ${{}_{(u,u)}}{\bf M}^{\prime}$ for $\mathbb{A}$-$C$-bimodule $\bf M$} and $\mathbb{A}_{\delta}$-$C$-bimodule $\bf M$$^{\prime}$, respectively.
Corollary 4.10.
Let $u$ be an element in $C_{P}(D)$. Then for any $\mathbb{A}$-module $M$, it is isomorphic to its twisted module ${{}_{u}}M$ as $\mathbb{A}$-modules. In particular, if $D$ is in the center $Z(P)$ of $P$, every $\mathbb{A}$-module $M$ is isomorphic to its twisted module ${{}_{v}}M$ as $\mathbb{A}$-modules for any $v\in P$.
Proof.
We adopt the notation in the paragraph above and the proof of Proposition 4.6. For the sake of clarity, we denote the element $m$ in $M$ by ${{}_{u}}m$ when we view it as an element in the twisted module ${{}_{u}}M$. We use the analogous notation for $\mathbb{A}_{\delta}$-modules.
[…]
${\bf f}(\mathfrak{a}_{\delta}\cdot({{}_{v}}m))=$~ $)\displaystyle{\bf f}(j\cdot({{}_{v}}(v^{-1}\mathfrak{a}_{\delta}v\cdot m)$))
[…]
$=$~ $\mathfrak{a}_{\delta}\cdot{\bf f}(j\cdot({{}_{v}}m)),$
which means that $\bf f$} is an isomorphism of $\mathbb{A}_{\delta}$-modules.
Next, we show that the element $a_{u,\delta}$ can be chosen such that it acts trivially on the subalgebra $\mathbb{B}$ through the conjugation action. Recall that $\hat{E}_{\mathfrak{h}}=\{\mathfrak{a}_{g}\mathfrak{a}_{\delta}\mid\tilde{g}\in E_{\mathfrak{h}},\mathfrak{a}_{\delta}\in(\mathbb{A}_{\delta}^{D})^{\times}\}$ is a subgroup of $N_{\mathbb{A}_{\delta}}(D)$ containing $(\mathbb{A}_{\delta}^{D})^{\times}$. Since $u$ belongs to $C_{P}(D)$, we have $\mathfrak{a}_{g}^{-1}{{}^{a_{u,\delta}}}\mathfrak{a}_{g}$ belonging to $(\mathbb{A}_{\delta}^{D})^{\times}$. In particular, we have ${{}^{a_{u,\delta}}}\hat{E}_{\mathfrak{h}}=\hat{E}_{\mathfrak{h}}$. Therefore, ${{}^{a_{u,\delta}}}\mathbb{M}/k^{\times}$ is also a complement of $j+J(\mathbb{A}_{\delta}^{D})$ in $\hat{E}_{\mathfrak{h}}/k^{\times}$. Then by [29, Lemma (45.6)], ${{}^{a_{u,\delta}}}\mathbb{M}$ is conjugate to $\mathbb{M}$ by an element in $j+J(\mathbb{A}_{\delta}^{D})$. We can adjust the element $a_{u,\delta}$ by multiplying this element in $j+J(\mathbb{A}_{\delta}^{D})$ such that ${{}^{a_{u,\delta}}}\mathbb{M}=\mathbb{M}$. Without loss of generality, we can set $\mathbb{M}=\{\mathfrak{a}_{g}\mu\mid\tilde{g}\in E_{\mathfrak{h}},\mu\in k^{\times}\}$. So for any $\tilde{g}\in E_{\mathfrak{h}}$, ${{}^{a_{u,\delta}}}\mathfrak{a}_{g}=\mu_{g}\mathfrak{a}_{g}$ for some $\mu_{g}\in k^{\times}$. Since $u$ is a $p$-element, there is some positive integer $a$ such that $a_{u,\delta}^{p^{a}}$ belongs to $(\mathbb{A}_{\delta}^{D})^{\times}$. Then $a_{u,\delta}^{p^{a}}=\mathfrak{a}_{\delta}\lambda$ for a unique element $\mathfrak{a}_{\delta}$ in $j+J(\mathbb{A}_{\delta}^{D})$ and a unique element $\lambda$ in $k^{\times}$. Hence,
${{}^{\mathfrak{a}_{\delta}}}\mathfrak{a}_{g}={{}^{a_{u,\delta}^{p^{a}}}}\mathfrak{a}_{g}=\mu_{g}^{p^{a}}\mathfrak{a}_{g}.$
Multiplying by $\mathfrak{a}_{g}^{-1}$ on the right yields an equation ${{}^{\mathfrak{a}_{\delta}}}\mathfrak{a}_{g}\mathfrak{a}_{g}^{-1}=\mu_{g}^{p^{a}}$. Note that the element in the left side belongs to $j+J(\mathbb{A}_{\delta}^{D})$ and $\mu_{g}^{p^{a}}$ is an element in $k^{\times}$. So $\mu_{g}$ has to be $1$. This completes the proof of the argument in this paragraph.
Moreover, by Corollary 4.8, there is an element $z_{u,\delta}$ in $C_{\mathbb{A}_{\delta}}(\mathbb{B})^{\times}$ such that ${{}^{a_{u,\delta}}}\mathfrak{c}_{\delta}={{}^{z_{u,\delta}}}\mathfrak{c}_{\delta}$ for any $\mathfrak{c}\in\mathbb{A}_{\delta}$. Then it is easy to check that for any $\mathbb{A}_{\delta}$-module $M_{\delta}$, it is isomorphic to its twisted module ${{}_{u}}M_{\delta}$. Applying this fact to the $\mathbb{A}$-module $jM$, we can get the following isomorphisms of $\mathbb{A}_{\delta}$-modules
$j({{}_{u}}M)\cong{{}_{u}}(jM)\cong jM.$
By Corollary 3.4, the $\mathbb{A}$-$\mathbb{A}_{\delta}$-bimodule $\mathbb{A}j$ induces a Morita equivalence between $\mathbb{A}$ and $\mathbb{A}_{\delta}$. Hence, we have ${{}_{u}}M$ isomorphic to $M$ as $\mathbb{A}$-modules. ∎
4.11.
For the remainder of this section we assume that the block $b$ is a hyperfocal abelian Frobenius block. We will give a more explicit description of the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $X$ in the decomposition (4.6.1) under this assumption. To do this, we need to investigate the Brauer quotients of $\mathbb{A}_{\delta}$ at all nontrivial subgroups of $D$. We start with a special case.
@@
Denote $N_{P}(Q)$ by $P_{Q}$. It is clear that $b_{Q}$ is still a block of $P_{Q}C_{G}(Q)$ over $\mathcal{O}$. We denote it by $c_{Q}$ when the block $b_{Q}$ is viewed as a block of $P_{Q}C_{G}(Q)$. By the assumption, $P_{Q}$ is a defect group of the block $c_{Q}$. Since the block $b_{Q}$ is nilpotent by Proposition 3.12, the block $c_{Q}$ of $P_{Q}C_{G}(Q)$ is also nilpotent. Since $P_{Q}$ contains $D$, there is a unique local pointe $\varepsilon$ of $P_{Q}$ on $\mbox{$\mathcal{O}$}Gb$ such that $(P_{Q})_{\varepsilon}\leq P_{\gamma}$ by [25, Proposition 4.2]. Now we can take an element $l$ of $\varepsilon$ such that $l$ is in $\mathbb{A}$ by [25, Proposition 2.8] and $lj=jl=j$.
By [29, Lemma (40.2)], $\mathrm{Br}_{Q}(l)$ belongs to a local point of $P_{Q}$ on $kC_{G}(Q)$. Since $P_{Q}$ is a defect group of the block $c_{Q}$, $\mathrm{Br}_{Q}(l)kP_{Q}C_{G}(Q)\mathrm{Br}_{Q}(l)$ is a source algebra of the block $\bar{c}_{Q}$ over $k$. Since the block $c_{Q}$ is nilpotent, by [16, Theorem 8.12.3], there is a an indecomposable endopermutation $kP_{Q}$-module $V_{P_{Q}}$ with vertex $P_{Q}$ such that, setting $\bar{T}_{Q}=\mathrm{End}_{k}(V_{P_{Q}})$, we have an isomorphism of $P_{Q}$-interior algebras
$\mathrm{Br}_{Q}(l)kP_{Q}C_{G}(Q)\mathrm{Br}_{Q}(l)\cong\bar{T}_{Q}\mathop{\otimes}\limits_{k}kP_{Q},$
@@
It is well-known that there is a unique local point, denoted by $\tilde{\delta}$, of $D$ on $\bar{T}_{Q}$. Take an element $f$ in $\tilde{\delta}$ and set $V_{Q}=fV_{P_{Q}}$. Then $V_{Q}$ is an indecomposable endopermutation $kD$-module with vertex $D$. Clearly, the action of $Q$ on $V_{Q}$ is also trivial. So $V_{Q}$ can be viewed as a $k(D/Q)$-module with vertex $D/Q$. Set $\bar{S}_{Q}=\mathrm{End}_{k}(V_{Q})\cong f\bar{T}_{Q}f$. By [16, Proposition 7.4.4], the $D$-interior algebras $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kC_{P}(Q)$ and $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ are both primitive. Note that $\mathrm{Br}_{Q}:\mathbb{A}^{D}\rightarrow\mathbb{A}(Q)^{D/Q}$ is surjective and $\mathrm{Br}_{D/Q}\circ\mathrm{Br}_{Q}$ is just the Brauer homomorphism $\mathrm{Br}_{D}$ from $\mathbb{A}^{D}$ to $\mathbb{A}(D)$. Then $\mathrm{Br}_{Q}(j)$ belongs to a local point of $D$ on $((l\mathbb{A}l)(Q))^{D}$. Comparing the isomorphism (4.12.3) with the isomorphism (4.12.4), we get $(lAl)(Q)=\mathop{\bigoplus}\limits_{uD\in C_{P}(Q)/D}(l\mathbb{A}l)(Q)\cdot u$. Therefore, by [25, Proposition 2.8], $\mathrm{Br}_{Q}(j)$ also belongs to a local point of $D$ on $((lAl)(Q))^{D}$. Note that by [22, 5.3], local points of $D$ on $(lAl)(Q)$ and $(l\mathbb{A}l)(Q)$ are both unique. Therefore, we can assume that the isomorphism (4.12.3) maps $\mathrm{Br}_{Q}(j)$ to $f\mathop{\otimes}\limits_{k}1_{C_{P}(Q)}$. Then the isomorphisms (4.12.1) and (4.12.2) are obtained.
Borrowing the notation from the paragraph 3.1, we have $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}j\mathbb{A}uj$ with $a_{u,\delta}\in A_{\delta}^{\times}\cap\mathbb{A}u$ satisfying $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ for any $w\in D$. So $A_{\delta}(Q)=\mathop{\bigoplus}\limits_{uD\in P/D}(j\mathbb{A}uj)(Q)$ and $(j\mathbb{A}uj)(Q)=\mathbb{A}_{\delta}(Q)\cdot\mathrm{Br}_{Q}(a_{u,\delta})$ when $u$ is in $C_{P}(Q)$. In particular, $\mathrm{dim}_{k}(\mathbb{A}_{\delta}(Q))=\mathrm{dim}_{k}((j\mathbb{A}uj)(Q))$ when $u$ is in $C_{P}(Q)$. By isomorphisms (4.12.1) and (4.12.2), we have $\mathrm{dim}_{k}(A_{\delta}(Q))=\frac{|C_{P}(Q)|}{|D|}\mathrm{dim}_{k}(\mathbb{A}_{\delta}(Q))$. We are done. ∎
By using Lemma 4.12, we can obtain the strucutre of $\mathbb{A}_{\delta}(Q)$ for arbitrary nontrivial subgroup $Q$ of $D$. And this general structure will help us construct a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ in the next section.
@@
Moreover, if $V_{Q}^{\prime}$ is another indecomposable endopermutation $k(D/Q)$-module with vertex $D/Q$ such that setting $\bar{S}_{Q}^{\prime}=\mathrm{End}_{k}(V_{Q}^{\prime})$, $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras, then $V_{Q}$ is isomorphic to $V_{Q}^{\prime}$ as $k(D/Q)$-modules.
@@
It is well-known that there is some $x\in G$ such that ${{}^{x}}(Q,\bar{b}_{Q})\leq(P,\bar{b}_{P})$ and $C_{P}({{}^{x}}Q)$ is a defect group of the block ${{}^{x}}b_{Q}$ of $C_{G}({{}^{x}}Q)$ over $\mathcal{O}$. Equivalently, $C_{P^{x}}(Q)$ is a defect group of the block $b_{Q}$. Since $D$ is abelian, by [33, Theomre 2], we can assume that $x$ is in $N_{G}(D_{\delta})$. Furthermore, since $N_{G}(D_{\delta})/C_{G}(D)=PC_{G}(D)/C_{G}(D)\cdot E_{\mathfrak{h}}$, we can assume that $xC_{G}(D)$ is in $E_{\mathfrak{h}}$. Then we have $\mathfrak{a}_{x}\in\mathbb{A}_{\delta}^{\times}$ such that $\mathfrak{a}_{x}(wj)\mathfrak{a}_{x}^{-1}=(xwx^{-1})j$ for any $w\in D$.
@@
Clearly, $A^{x}$ is also a source algebra of the block $b$ determined by the defect pointed group $(P_{\gamma})^{x}$. Since $x$ is in $N_{G}(D_{\delta})$, we have $\mathbb{A}^{x}$ is a hyperfocal subalgebra of $A^{x}$ with hyperfocal subgroup $D$. The conjugation action of $x$ induces an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{{{}^{x}}Q}\longrightarrow((\mathbb{A}^{x})_{\delta^{x}})^{Q}$. At the same time, the conjugation action of $\mathfrak{a}_{x}$ induces an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{Q}\longrightarrow\mathbb{A}_{\delta}^{{{}^{x}}Q}$. Composing these two isomorphisms together, we get an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{Q}\longrightarrow(\mathbb{A}^{x})_{\delta^{x}}^{Q}$, which is in fact a $D$-interior algebra isomorphism. By Lemma 4.12, we get this general structure of $\mathbb{A}_{\delta}(Q)$.
@@
Let $V_{Q}^{\prime}$ be another indecomposable endopermutation $k(D/Q)$-module with vertex $D/Q$ fulfilling the property in this lemma. Then we have $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Tensoring with $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}$- on the both sides of this isomorphism yields a $D$-interior algebra isomorphism $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$. By [22, 5.3], local points of $D$ on $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime}$ and $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ are both unique. We take two elements $e^{\prime}$ and $e$ belonging to these two unqiue local points, respectively. Then by [16, Theorem 7.4.2], $e(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD)e$ is isomorphic to $kD$ as $D$-interior algebras. It is clear that $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}$ is also a primitive Dade $D$-interior algebra. By [16, Proposition 7.4.4], we have $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}\mathop{\otimes}\limits_{k}kD$ is a primitive $D$-interior algebra. Hence, we have $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}\mathop{\otimes}\limits_{k}kD\cong kD$ as $D$-interior algebras. This implies that $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}$ is a trivial $D$-interior algebra. Then the elements in the Dade group (see [16, Definition 7.3.11]) of $D$ corresponding to $V_{Q}$ and $V_{Q}^{\prime}$ are inverse to each other. In particular, $V_{Q}$ is isomorphic to $V_{Q}^{\prime}$ as $k(D/Q)$-modules. ∎
Remark 4.14.
With the notation of Proposition 4.13, $A_{\delta}(Q)$ may not be isomorphic to $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kC_{P}(Q)$ in general. It behaves different from the special case in Lemma 4.12. For example, let $G$ be the symmetric group ${\rm S}_{4}$ on $4$ letters and $p$ equal to $2$. Let $b$ be the principal block of $G$ over $\mathcal{O}$. In fact, $b$ is just the unit element $1_{\mathcal{O}G}$ and the source algebra $A$ is equal to the group algebra $\mbox{$\mathcal{O}$}G$. In this case, the hyperfocal subgroup $D$ is a Klein four group and the hyperfocal subalgebra $\mathbb{A}$ is just the group algebra $\mbox{$\mathcal{O}$}{\rm A}_{4}$, where ${\rm A}_{4}$ is the alternating group. Set $P$ to be $D\rtimes(12)$. Let $Q$ be the subgroup $\{(1),(13)(24)\}$ of $D$. Then $C_{P}(Q)$ is just equal to $D$ and $C_{G}(Q)$ is equal to another Sylow $2$-subgroup $P^{\prime}=D\rtimes(13)$. In this case, $A_{\delta}$ is isomorphic to $kC_{G}(Q)$, which is $kP^{\prime}$.
Due to this, in general $(j\mathbb{A}vj)(Q)$ may be not zero when $v$ does not belong to $C_{P}(Q)$ since in the concrete example above $(j\mathbb{A}vj)(Q)$ is equal to $(kD)(13)$ when $v$ equals $(12)$ not in $C_{P}(Q)$.
4.15.
@@
Let $Q$ be a subgroup of $D$ and $\varphi$ a homomorphism from $Q$ to $D$. We denote the subgroup $\{(u,\varphi(u))\mid u\in Q\}$ of $D\times D$ by $\Delta_{\varphi}(Q)$. For any $g\in G$ with $Q^{g}\leq D$, we denote by $\varphi_{g}$ the homomorphism from $Q$ to $D$ induced by the conjugation action of $g$. Now we can refine the $\mbox{$\mathcal{O}$}(D\times D)$-structure of $\mathbb{A}_{\delta}$ in the case where the block $b$ is a hyperfocal abelian Frobenius block as follows.